### v1

@@ V Optimal Wideband Hybrid Beamformers: A Cost-Performance Trade-off Analysis
Figure 5: Optimal structure for beam squint mitigation in wideband hybrid beamformers considering the cost-performance trade-off. The normalized performance and cost benefits in each of the wideband and ultra-wideband scenarios are also visualized.
VI Conclusions
@@
In this work, we presented a comparative evaluation of Non-TTD, Sparse-TTD, and Full-TTD hybrid beamforming architectures for mitigating beam squint in wideband and ultra-wideband communication systems. For each architecture, we examined the underlying technological principles, implementation challenges, and recent advancements reported in the literature. A key insight from our analysis is the often-overlooked practical limitation imposed by wideband phase shifters, which significantly affects the cost, complexity, and performance of both Non-TTD and Sparse-TTD architectures. We discussed how these components constrain the practical efficiency of such designs, particularly in ultra-wideband scenarios. In this context, our study concludes that Full-TTD-based beamformers offer the most efficient solution for ultra-wideband systems, not only in terms of performance but also in terms of cost, since they eliminate the need for wideband phase shifters. Conversely, for moderate-wideband systems, Sparse-TTD beamformers or Non-TTD-based structures such as Non-TTD-WBBG may be preferable, depending on whether performance or cost is prioritized.

### v2

@@ VI Optimal Wideband Hybrid Beamformers: A Cost-Performance Trade-off Analysis
Figure 5: Beam squint mitigation in wideband hybrid beamformers considering the cost-performance trade-off. The normalized performance and cost benefits in each of the wideband and ultra-wideband scenarios are also visualized.
VII Conclusions
@@
In this work, we presented a comparative evaluation of Non-TTD, Sparse-TTD, and Full-TTD hybrid beamforming architectures for mitigating beam squint in wideband and ultra-wideband communication systems, examining for each the underlying technological principles, implementation challenges, and recent advancements reported in the literature. A key insight from our analysis is the often-overlooked practical limitation imposed by wideband PSs, which significantly affects the cost, complexity, and performance of both Non-TTD and Sparse-TTD architectures, constraining their practical efficiency particularly in ultra-wideband scenarios. Accordingly, our study concludes that Full-TTD-based beamformers potentially offer the most efficient solution for ultra-wideband systems not only in terms of performance but also in terms of cost, since they eliminate the need for wideband PSs.